OpenStudy (anonymous):

3sq root 72/4 sq root 98/2

OpenStudy (mertsj):

\[3\sqrt{\frac{72}{4}}\sqrt{\frac{98}{2}}=3\sqrt{18}\sqrt{49}=3\sqrt{9(2)}(7)=21(3)\sqrt{2}=63\sqrt{2}\]